OpenStudy (anonymous):

Find the sum of the arithmetic series: 6 + 10 + 14 + ... + 86

OpenStudy (anonymous):

plus 4 each time

OpenStudy (calculusxy):

The answer should be 996

hartnn (hartnn):

same type as last one.... did u try first ?

OpenStudy (anonymous):

a=6,l=86 find n by l=a+(n-1)d then Sn=n(a+l)/2

OpenStudy (anonymous):

so 21=n?

OpenStudy (anonymous):

yea,correct

OpenStudy (anonymous):

so then what

OpenStudy (anonymous):

Sn=21(6+86)/2

OpenStudy (anonymous):

966???

OpenStudy (anonymous):

correct

hartnn (hartnn):

yeah 966 is correct, now can you solve similar type of questions on your own ?

OpenStudy (anonymous):

I think so...